OpenStudy (anonymous):

can someone check if my answer is correct, heres the problem, two identical positive charges has been separated by a 2.6 x 10^-2m, being released from rest. immediaately after the release, particle 1 has an accelelration a1 whose magnitude is 4.6 x 10^3m/s^2, while particle 2 has an acceleration a2 whose magintude is 8.5 x 10^3 m/s^2. find the unknown charge on each paticle and the mass of particle 2

OpenStudy (anonymous):

F = ma F = (6x10^-6)(4.6 x 10^3) F = 0.0276 N F = kq1q2/r^2 o.o276 = 9.0 x 10^9(x^2)/ (2.6x10^-2)^2 x = 5. 078 x 10^-8 mass1 m = f/a = 0.0276/8.5x10-3 m= 3.247x10^-6

OpenStudy (anonymous):

@irishboy123

OpenStudy (anonymous):

@ElonaSushchik

OpenStudy (elonasushchik):

I think you are correct!

OpenStudy (anonymous):

how to create the fbd ?

OpenStudy (anonymous):

@BlossomCake @january123

OpenStudy (irishboy123):

makes no sense to me your first line: F = ma what are F m and a in that one? i think some information is missing.

OpenStudy (irishboy123):

have you progressed this?